# Programmer Visual Attention During Context-Aware Code Summarization

### v1

@@ Abstract
Programmer attention represents the visual focus of programmers on parts of the source code in pursuit of programming tasks. The focus of current research in modeling this programmer attention has been on using mouse cursors, keystrokes, or eye-tracking equipment to map areas in a snippet of code. These approaches have traditionally only mapped attention for a single method. However, programming tasks such as source code summarization, require programmers to use contextual knowledge that can only be found in other parts of the project. To address this knowledge gap, we conducted an in-depth human study with XY Java programmers, where each programmer generated summaries for 40 methods from five large Java projects over five one-hour sessions. We used eye-tracking equipment to map the visual attention of programmers while they wrote the summaries. We also rate the quality of each summary. We found eye-gaze patterns and metrics that define common behaviors between programmer attention during context-aware code summarization. Specifically, we found that programmers need to read significantly (p<0.01) fewer words and make significantly fewer revisits to words (p<0.03) as they summarize more methods during a session, while maintaining the quality of summaries. We also found that the amount of source code a participant looks at correlates with a higher quality summary, but this trend follows a bell-shaped curve, such that after a threshold reading more source code leads to a significant decrease (p<0.01) in the quality of summaries. We also gathered insight into the type of methods in the project that provide the most contextual information for code summarization based on programmer attention. Specifically, we observed that programmers spent a majority of their time looking at methods inside the same class as the target method to be summarized. Surprisingly, we found that programmers spent significantly less time looking at methods in the call graph of the target method. We discuss how our empirical observations may aid future studies towards modeling programmer attention and improving context-aware automatic source code summarization.
@@ 1 Introduction
However, these summaries do not necessarily reflect real-world scenario, as they are not “context-aware”. A summary is context-aware when it considers the context, i.e., information from outside the subroutine. Although there are several types of context, for the purposes of this study, context is source code from the project surrounding the subroutine. Context-aware summaries are important because a subroutine usually does not exist in a vacuum; the manner in which a subroutine is used depends on the context. Therefore project-level information in critical to writing a context-aware summary. Over the last five years, a few studies have introduced context to code summarization models, such as by adding source code from part of the project  [2, 3] or the call-graph [4]. It is cost-prohibitive to feed the entire project to these language models as a longer input requires more computational resources. There is a knowledge gap in related work on how human programmers navigate through the context in a project to infer the most important information. That is the core inspiration behind this study. We designed this study to map the programmers’ visual attention strategy and find out what in the context they find most important, so that we may use that information for future studies in automated source code summarization.
@@
The visual attention strategies of a programmer are an important indication of their mental model when performing programming tasks [5, 6, 7]. Visual attention strategies refer to the selective focus humans apply to some parts of the visual stimulus (i.e., the code in the case of programmers). This selection of parts of the code is informed by the mental model of the programmer that evolves as they retain more information [8]. These visual attention strategies can be inferred using eye-tracking data, specifically, gaze patterns of programmers. Eye-trackers have been used in fields such as computer vision, psychology, and medical sciences to create models of human behavior and mental processes for decades [9, 10, 11]. Although models of visual attention strategies are interesting on their own, they can used to learn domain-specific knowledge such as those programmers use while completing SE tasks.
@@
To that end, recent work in SE towards modeling programmer attention has mainly used proxies such as mouse cursor movements and clicks [12]. Studies to model visual attention directly using eye-trackers have been limited to a small snippet of code [13, 6, 14]. Recently, a few studies have also proposed approaches for automatic inference of visual attention [15, 16, 17]. The scope of most of these studies is limited to one subroutine. We build upon these works to study eye-tracking and visual attention patterns at the project-level. Compared to those studies, our study design is challenging due to difficulty controlling a study environment where the participants have access to the entire project. Analyzing the data at the project level also poses challenges as not all code in the project is useful or even executed. However, project-level comprehension is a real world scenario in which programmers complete SE tasks such as code summarization. To the best of our knowledge, we conducted the first comprehensive human study to map these attention patterns in the context of the project.
@@
In this paper, we present an eye-tracking study designed with the goal of analyzing visual attention strategies of programmers while they navigate related methods for the task of source code summarization. We recruited 10 Java programmers, each tasked with completing five sessions. In each session, we asked programmers to write context-aware summaries for 8 subroutines of a Java project. Then, we manually rated the summaries for accuracy, completeness, conciseness, and clarity using two graders, who are among the authors. These ratings help us analyze the effect of various programmer gaze patterns on the quality of summaries. Finally, we categorized all the code in the project into categories of code context. We then identified the type of context to which the programmers gave the highest attention in pursuit of the summarization tasks. Although knowledge of programmer visual attention strategies is intellectually interesting, we posit that models of programmer attention and identification of important parts of context have practical applications towards improving models for automatic source code summarization.
@@
We found that: 1) participants looked at fewer words and revisited words fewer times as they processed more methods for a given project; 2) the quality of these summaries did not suffer or decrease with any statistical significance as the participants processed more tasks; 3) the quality of summaries is related to the amount of context analyzed by the participants, like a bell-curve, such that the quality increases as more context is analyzed by the participants upto a certain threshold and then the quality decreases; 4) participants spent the most time looking at class methods, followed by class declarations and call graphs; 5) the type of context participants focused on is generally not affected by the project, session order, method order, or participant.
@@
    We perform an eye-tracking study involving more than 60 hours of participant effort by 10 Java programmers. Each programmer completed 40 context-aware code summarization tasks.
@@ 2 Background and Related Work
2.1 Eye-tracking in Software Engineering
@@
Eye-tracking has been used in SE research for nearly two decades to gain insight into programmer behavior and human factors in programming [7]. Recent evolution of eye-trackers from expensive and difficult-to-operate lab equipment to simple monitor-mounted devices usable in development environment has accelerated their use in SE research [18, 16]. Eye-tracking studies in SE can mainly be divided into three areas of interest:
@@
Code comprehension is an area of interest that is primarily concerned with the mapping of programmer mental model during comprehension of a snippet of code or small program. In 2006, Aschwanden and Crosby [19] presented gaze patterns of programmers during the comprehension of a small snippet of code containing a loop and mathematical equations. They formalized the process of understanding programmer mental processes using eye-tracking during program comprehension. Also in 2006, Bednarik and Tukiainen [13] introduced a framework for analyzing eye gaze patterns, in particular scanpaths, of participants tasked with comprehension of two small programs. They found that the mental model of a programmer is refined as they progress through the task, marked by a decrease in attention switching. In 2014, Rodeghero et al. [20] presented an analysis of gaze patterns of programmers tasked with summarization of source code. They used the findings of that study to improve automatic source code summarization. Similarly, Karas et al. [14] presented and compared visual attention strategies of programmers for two tasks: reading of source code summaries for assessment and generating their own source code summaries. They found that programmers tend to look at specific words in the method in specific order defined by the syntactic nature of the code. Their work serves as an inspiration for this study, but these studies only considered method-level information. In this study, we focus our analysis on the project-level information.
@@
UML Diagrams are an area of interest concerned with understanding how programmers use UML diagrams during programming tasks. In 2006, Guéhéneuc [21] presented two case studies on the importance of UML diagrams in program comprehension by following the gaze patterns of programmers presented with a comprehension task. In 2010, Sharif and Maltec [22] showed that eye gaze patterns of programmers suggest better program comprehension when presented with particular UML layouts and clustering patterns. A caveat of studies on UML diagrams is that the results are specific to the visualization tools and programmers were not presented source code or asked to do a programming task [23, 24]. Therefore, the observations from these studies do not necessarily represent visual attention during code comprehension.
@@
Code Debugging studies are concerned with mapping programmer attention and mental model during the task of code debugging [7]. In 2008, Bednarik and Tukiainen [8] presented an analysis of gaze patterns of programmers with varying levels of experience during debugging tasks. They found that more experienced programmers tend to use context-switching quicker and more often that novice programmers. In 2014, Turner et al. [25] presented results of a study performed exclusively with students given two tasks: 1) over-viewing code to answer questions, and 2) debugging code. They found that gaze patterns are significantly different between the two tasks, suggesting eye-patterns of code comprehension are goal-dependent. This is important because it highlights the need for in-depth studies like ours that specifically ask the programmer to summarize code using context.
@@
Our study is most closely related to code comprehension as we present participants with raw source code and specifically task them with summarization using context. The novelty of our study lies with the project-level eye-tracking and data analysis.
@@ 2.2 Source Code Summarization
In 2023, Bansal et al. [4] introduced a Graph Neural Network (GNN) based encoder to represent the call graph extracted from the project centered at the target subroutine. Recently, two studies used human visual attention to improve automatic source code summarization by using synthesized human attention data given limited eye-tracking data [15, 17]. However, the short summaries generated by these approaches may not represent the full potential of contextual information. Therefore, in this study we ask participants to write longer summaries that include specific contextual information in them.
@@ 3.1 Study Scenario
We designed this study to emulate a specific scenario in code comprehension, specifically context-aware source code summarization. We asked participants to retrospectively write context-aware descriptions of a target method. We provide participants access to the entire Java code from the project that contains the target method. The project is devoid of any documentation, including header and inline comments (See on details of the projects). We ask the programmer to read and understand the target method, as well as the surrounding code in the context, and to write a three-sentence context-aware summary with the following instructions:
@@
We asked each programmer to write summaries for 8 methods per project, found to be the ideal number for 1 hour session during our pilot studies. We chose 5 projects extracted from Github for this study, and displayed only the raw Java code. We ask every programmer to process each project in a separate eye-tracking session. We set up the study entirely in Eclipse IDE, a popular IDE for open source Java developers to emulate an OSS development environment. See Section 3 for complete study procedure, interface and other technical details.
3.2 Research Question
We answer three main research questions:
@@ 3.3 Interface
We use iTrace suite [38, 39] for our eye-tracking study. The suite includes three tools: 1)iTrace core [40] that manages the eye-tracking, screen recording, keyboard, and mouse recording; 2) iTrace Eclipse plugin [37] that connects to the core and maps the gaze data to specific elements of the IDE and the text inside files; 3) textbfiTrace toolkit [41] which is a post-processing tool to isolate gaze patterns. We used this suite because it is defined to be used over an IDE and can record eye-tracking data over the entire code project.
@@ 3.5 Study Procedure
  • Before the participant arrives, we launched both the Tobii eye-tracker manager and iTrace core program we set up the session and participant details.
@@
  • We ask the participant to perform two calibrations before starting the first task as recommended by itrace developers [40]. First, Tobii eye-tracker manager and then the iTrace core calibration. After calibration, we ask the participant to start a tracking session on the iTrace core window and begin the task inside the Eclipse IDE.
  • To complete a task, we ask the participant to navigate to a file and line number and find the method. Then, the participant is free to explore any number of files in the project. After the completion of one task, we designed the text file to remind the participant to stop tracking and start a new tracking session for the next task. This allows each task to have its own eye-tracking data and screen recording file.
  • The study administrator helps re-calibrate with both Tobii eye-tracker manager and iTrace halfway through the session, after task 4, and before starting the new tracking session for task 5. We also prompt re-calibration if the participant needs a break or leaves the eye-tracking room at any point during the session.
@@ 3.6 Java Projects
We scraped Github for repositories with atleast 90% of the project code base containing Java source code. We also filtered for projects that depend heavily on other projects or codebases. Finally we picked the following five projects:
  • Scrimage is an immutable, functional, and performant JVM library for manipulation of images. The aim of this library is to provide a simple and concise way to do common image operations, such as resizing, filter, and converting between formats.
@@
Note, for each project above we chose 8 specific methods. To ensure that the methods can be described using relevant context, we prioritized methods that are called by at least one other method in the project. The order of methods 2 through 8 is randomized. We do this to answer RQ1 by minimizing the effect of any one method on the observations, i.e., a particularly tough or easy method would not skew the results as it would appear at a different position for each participant. Method 1 is fixed for each session to start each participant off with the same baseline.
3.7 Eye-Tracker
The eye-tracker used for this study was the Tobii Pro Fusion at 120Hz mounted at the bottom of a Spectre 24-inch 1920x1080 resolution monitor at 60Hz screen refresh rate. The eye-tracker has an accuracy of 0.03 and precision of 0.04 in optimal conditions [43]. Note, although all sessions were conducted in the same lab with the same equipment, we could not ensure optimal conditions were met for all participants due to factors such as ambient light, natural light and corrective lenses.

### v2

@@ Abstract
Programmer attention represents the visual focus of programmers on parts of the source code in pursuit of programming tasks. The focus of current research in modeling this programmer attention has been on using mouse cursors, keystrokes, or eye tracking equipment to map areas in a snippet of code. These approaches have traditionally only mapped attention for a single method. However, there is a knowledge gap in the literature because programming tasks such as source code summarization require programmers to use contextual knowledge that can only be found in other parts of the project, not only in a single method. To address this knowledge gap, we conducted an in-depth human study with 10 Java programmers, where each programmer generated summaries for 40 methods from five large Java projects over five one-hour sessions. We used eye tracking equipment to map the visual attention of programmers while they wrote the summaries. We also rate the quality of each summary. We found eye-gaze patterns and metrics that define common behaviors between programmer attention during context-aware code summarization. Specifically, we found that programmers need to read up to 35% fewer words (p<0.01) over the whole session, and revisit 13% fewer words (p<0.03) as they summarize each method during a session, while maintaining the quality of summaries. We also found that the amount of source code a participant looks at correlates with a higher quality summary, but this trend follows a bell-shaped curve, such that after a threshold reading more source code leads to a significant decrease (p<0.01) in the quality of summaries. We also gathered insight into the type of methods in the project that provide the most contextual information for code summarization based on programmer attention. Specifically, we observed that programmers spent a majority of their time looking at methods inside the same class as the target method to be summarized. Surprisingly, we found that programmers spent significantly less time looking at methods in the call graph of the target method. We discuss how our empirical observations may aid future studies towards modeling programmer attention and improving context-aware automatic source code summarization.
@@ 1 Introduction
However, these summaries do not necessarily reflect real-world scenario, as they are not “context-aware”. A summary is context-aware when it considers the context, i.e., other code in the project related the subroutine either required to run it or understand it. [2, 3] Although there are several types of context, for the purposes of this study, context is source code from the project related to the subroutine. Context-aware summaries are important because a subroutine usually does not exist in a vacuum; the manner in which a subroutine is used depends on the context. Therefore project-level information in critical to writing a context-aware summary. Over the last five years, a few studies have introduced context to code summarization models, such as by adding source code from part of the project  [4, 5] or the call-graph [6]. It is cost-prohibitive to feed the entire project to these language models as a longer input requires more computational resources. There is a knowledge gap in related work on how human programmers navigate through the context in a project to infer the most important information. That is the core inspiration behind this study. We designed this study to discover what programmers’ visual attention is during code summarization and to discover what contextual information they find most important, so that we may use that information for future studies in automated source code summarization.
@@
The visual attention of a programmer is an important indication of their mental model when performing programming tasks [7, 8, 9]. Visual attention strategies refer to the selective focus humans apply to some parts of the visual stimulus (i.e., the code in the case of programmers). This selection of parts of the code is informed by the mental model of the programmer that evolves as they retain more information  [10]. These visual attention strategies can be inferred using eye tracking data, specifically, gaze patterns of programmers. eye trackers have been used in fields such as computer vision, psychology, and medical sciences to create models of human behavior and mental processes for decades [11, 12, 13]. Although models of visual attention strategies are interesting on their own, they can used to learn domain-specific knowledge such as those programmers use while completing SE tasks.
@@
To that end, recent work in SE towards modeling programmer attention has mainly used proxies such as mouse cursor movements and clicks [14]. Studies to model visual attention directly using eye trackers have been limited to a small snippet of code [15, 8, 16]. Recently, a few studies have also proposed approaches for automatic inference of visual attention [17, 18, 19]. The scope of most of these studies is limited to one subroutine. We build upon these works to study eye tracking and visual attention patterns at the project-level. Compared to those studies, our study design is challenging due to difficulty controlling a study environment where the participants have access to the entire project. Analyzing the data at the project level also poses challenges as not all code in the project is useful or even executed. However, project-level comprehension is a real world scenario in which programmers complete SE tasks such as code summarization. To the best of our knowledge, we conducted the first comprehensive human study to map these attention patterns in the context of the project.
@@
In this paper, we present an eye tracking study designed with the goal of analyzing visual attention strategies of programmers while they navigate related methods for the task of source code summarization. We recruited 10 Java programmers, each tasked with completing five sessions. In each session, we asked programmers to write context-aware summaries for 8 subroutines of a Java project. Then, we manually rated the summaries for accuracy, completeness, conciseness, and clarity using two graders, who are among the authors. These ratings help us analyze the effect of various programmer gaze patterns on the quality of summaries. Finally, we categorized all the code in the project into categories of code context. We then identified the type of context to which the programmers gave the highest attention in pursuit of the summarization tasks. Although knowledge of programmer visual attention strategies is intellectually interesting, we posit that models of programmer attention and identification of important parts of context have practical applications towards improving models for automatic source code summarization.
@@
We found that: 1) participants looked at fewer words and revisited words fewer times as they processed more methods for a given project; 2) the quality of these summaries did not suffer or decrease with any statistical significance as the participants processed more tasks; 3) the quality of summaries is related to the amount of context analyzed by the participants, like a bell-curve, such that the quality increases as more context is analyzed by the participants up to a certain threshold and then the quality decreases; 4) participants spent the most time looking at class and instance methods, followed by class declarations and call graphs; 5) the type of context participants focused on–such as class and instance methods, declarations, the call graph, or other areas of the project–is generally not affected by the project, session order, method order, or participant.
@@
    We perform an eye tracking study involving more than 60 hours of participant effort by 10 Java programmers. Each programmer completed 40 context-aware code summarization tasks.
@@ 2 Background and Related Work
2.1 Eye tracking in Software Engineering
@@
Eye tracking has been used in SE research for nearly two decades to gain insight into programmer behavior and human factors in programming [9]. Recent evolution of eye trackers from expensive and difficult-to-operate lab equipment to simple monitor-mounted devices usable in development environment has accelerated their use in SE research [20, 18]. Eye tracking studies in SE can mainly be divided into three areas of interest:
@@
Code Comprehension is an area of interest that is primarily concerned with the mapping of programmer mental model during comprehension of a snippet of code or small program. In 2006, Aschwanden and Crosby [21] presented gaze patterns of programmers during the comprehension of a small snippet of code containing a loop and mathematical equations. They formalized the process of understanding programmer mental processes using eye tracking during program comprehension. Also in 2006, Bednarik and Tukiainen [15] introduced a framework for analyzing eye gaze patterns, in particular scanpaths, of participants tasked with comprehension of two small programs. They found that the mental model of a programmer is refined as they progress through the task, marked by a decrease in attention switching. In 2014, Rodeghero et al. [22] presented an analysis of gaze patterns of programmers tasked with summarization of source code. They used the findings of that study to improve automatic source code summarization, showing how code summarization approaches of the time could be improved by making them focus on the same things as people do. Abid et al. [23] replicated and expanded the Rodeghero study [22] in 2019 with project context using a similar interface, and showed that allowing the participants to work in a more realistic environment results in a more generalized understanding of developer behavior.
Similarly, in 2024, Karas et al. [16] presented and compared visual attention strategies of programmers for two tasks: reading of source code summaries for assessment and generating their own source code summaries. They found that programmers tend to look at specific words in the method in specific order defined by the syntactic nature of the code. Their work serves as an inspiration for this study, however these studies only considered method-level information. In this study, we focus our analysis on the project-level information.
@@
UML Diagrams are an area of interest concerned with understanding how programmers use UML diagrams during programming tasks. In 2006, Guéhéneuc [24] presented two case studies on the importance of UML diagrams in program comprehension by following the gaze patterns of programmers presented with a comprehension task. In 2010, Sharif and Maletic [25] showed that eye gaze patterns of programmers suggest better program comprehension when presented with particular UML layouts and clustering patterns. A caveat of studies on UML diagrams is that the results are specific to the visualization tools and programmers were not presented source code or asked to do a programming task [26, 27]. Therefore, the observations from these studies do not necessarily represent visual attention during code comprehension.
@@
Code Debugging studies are concerned with mapping programmer attention and mental model during the task of code debugging [9, 28]. In 2008, Bednarik and Tukiainen [10] presented an analysis of gaze patterns of programmers with varying levels of experience during debugging tasks. They found that more experienced programmers tend to use context-switching quicker and more often that novice programmers. In 2014, Turner et al. [29] presented results of a study performed exclusively with students given two tasks: 1) over-viewing code to answer questions, and 2) debugging code. They found that gaze patterns are significantly different between the two tasks, suggesting eye-patterns of code comprehension are goal-dependent. In 2020, Sharafi et al. [30] presented a study that quantifiably indicates that developers generally follow a three-phase model of finding, learning, and editing. They found that developers spend more time searching code towards the beginning of the task, while what they focused on changed over time as they understood the program context. This is important because it highlights the need for in-depth studies like ours that specifically ask the programmer to summarize code using context, as well as finding out what parts of the code are more important to the programmer over time.
@@
Our study is most closely related to code comprehension as we present participants with raw source code and specifically task them with summarization using context. The novelty of our study lies with the project-level eye tracking and data analysis.
@@ 2.2 Source Code Summarization
In 2023, Bansal et al. [6] introduced a Graph Neural Network (GNN) based encoder to represent the call graph extracted from the project centered at the target subroutine. Recently, two studies used human visual attention to improve automatic source code summarization by using synthesized human attention data given limited eye tracking data [17, 19]. However, the short summaries generated by these approaches may not represent the full potential of contextual information. Therefore, in this study we ask participants to write longer summaries that include specific contextual information in them.
@@ 3.1 Study Scenario
We design this study to emulate a specific scenario in code comprehension, specifically context-aware source code summarization. We ask participants to retrospectively write context-aware descriptions of a target method. We provide participants access to the entire Java code from the project that contains the target method. The project is devoid of any documentation, including header and inline comments (See 3.6 on details of the projects). We ask the programmer to read and understand the target method, as well as the surrounding code in the context, and to write a three-sentence context-aware summary with the following instructions:
@@
We ask each programmer to write summaries for 8 methods per project, found to be the ideal number for 1 hour session during our pilot studies. We choose 5 projects extracted from Github for this study, and displayed only the raw Java code. We ask every programmer to process each project in a separate eye tracking session. We set up the study in the Eclipse IDE.
3.2 Research Questions
We answer three research questions:
@@ 3.3 Interface
We use iTrace suite [43, 44] for our eye tracking study. The suite includes three tools: 1) iTrace core [45] that manages the eye tracking, screen recording, keyboard, and mouse recording; 2) iTrace Eclipse plugin [46] that connects to the core and maps the gaze data to specific elements of the IDE and the text inside files; 3) iTrace toolkit [42] which is a post-processing tool to isolate gaze patterns. We use this suite because it can be used in an IDE and can record eye tracking data over the entire code project.
@@ 3.5 Study Procedure
  • Before the participant arrives, we launch both the Tobii eye tracker manager and iTrace core program. We then set up the session and participant details.
@@
  • We ask the participant to perform two calibrations before starting the first task as recommended by iTrace developers [45]. First, Tobii eye tracker manager and then the iTrace core calibration. After calibration, we ask the participant to start a tracking session on the iTrace core window and begin the task inside the Eclipse IDE.
  • To complete a task, we ask the participant to navigate to a file and line number and find the method. Then, the participant is free to explore any number of files in the project. After the completion of one task, we designed the text file to remind the participant to stop tracking and start a new tracking session for the next task. This allows each task to have its own eye tracking data and screen recording file.
  • The study administrator helps re-calibrate with both Tobii eye tracker manager and iTrace halfway through the session, after task 4, and before starting the new tracking session for task 5. We also prompt re-calibration if the participant needs a break or leaves the eye tracking room at any point during the session.
TABLE I: Description of Java Projects
Name Number of Files Number of Classes Lines of Code Age of Project (years) Domain
Scrimage 122 111 5981 11 Image Editing
MLTK 158 152 18816 2 Machine Learning
OpenAudible 107 80 13379 6 Multimedia
MALLET 619 767 77407 16 NLP
FreeCol 804 889 128049 20 Entertainment
@@ 3.6 Java Projects
We scraped Github for repositories with at least 90% of the project code base containing Java source code. We also excluded projects that depend heavily on other projects or codebases. Finally we picked the following five projects:
  • Scrimage is an immutable, functional, and performant Java library for manipulation of images. The aim of this library is to provide a simple and concise way to do common image operations, such as resizing, filter, and converting between formats.
@@
More details about the projects can be found in Table I. Notably, the projects cover a wide range of project domains, an 18 year age range, and between 5-128k lines of code. For each project above we chose 8 specific methods. To ensure that the methods can be described using relevant context, we avoided trivial methods such as getters, setters, or similar self-evident functions, while prioritizing methods that are called at least twice (for multiple direct contextual sources) or perform some functionality in the project discernible in the context of the project. The order of methods 2 through 8 is randomized. We do this to answer RQ1 by minimizing the effect of any one method on the observations, i.e., a particularly tough or easy method would not skew the results as it would appear at a different position for each participant. Method 1 is fixed for each session to start each participant off with the same baseline.
3.7 Eye tracker
The eye tracker used for this study was the Tobii Pro Fusion at 120Hz mounted at the bottom of a Spectre 24-inch 1920x1080 resolution monitor at 60Hz screen refresh rate. The eye tracker has an accuracy of 0.03 and precision of 0.04 in optimal conditions [48]. Note, although all sessions were conducted in the same lab with the same equipment, we could not ensure optimal conditions were met for all participants due to factors such as ambient light, natural light and corrective lenses.